# Neural Stereo Video Compression with Hybrid Disparity Compensation

### v1

@@ Abstract
1 Introduction
As stereoscopic cameras become increasingly common in newly launched smart products, such as virtual reality headsets and autonomous vehicles, the demand for efficient storage and transmission of large volumes of stereo video is growing. Stereo video compression (SVC) [55, 53, 25, 11] has recently gained significant attention for its ability to encode stereo videos into compact bitstreams, reducing both storage and transmission costs.
A key challenge in this task is reducing cross-view redundancy, and disparity compensation is central to this effort. By aligning stereo views, disparity compensation enables the extraction of missing details and the reuse of shared information, which enhances reconstruction quality and improves compression efficiency. Traditional stereo video codecs [53, 55] rely on handcrafted disparity compensation methods, but these approaches are limited by the constraints of manually designed structures. The recent success of deep neural networks (DNNs) in stereo visual data processing [49, 20, 29, 57, 65, 56, 15] has paved the way for neural-based disparity compensation mechanisms. These modern techniques offer a more flexible and accurate solution for reducing cross-view redundancy in stereo video compression.
@@ 1 Introduction
We conduct extensive experiments to evaluate our framework on widely used automotive stereo video benchmarks, including KITTI 2012 [19] and KITTI 2015 [46]. In contrast to previous NSVC methods [11, 25]—which have been evaluated solely on automotive scenes—we also test our approach on general multi-view scenes using the Nagoya dataset [3], a standard stereo video benchmark in MV-HEVC [55, 53]. The results indicate that our framework outperforms both neural and traditional SVC methods in both automotive and general scenes. Our contributions can be summarized as follows:
@@
  • We propose a hybrid disparity compensation (HDC) strategy that leverages explicit pixel displacement as a robust disparity prior to streamline optimization while incorporating implicit cross-attention mechanisms for improved compensation performance.
@@
  • We develop an end-to-end optimized NSVC framework by integrating our HDC into key components, including cross-view feature extraction and reconstruction (HDC-FER) and cross-view entropy modeling (HDC-EM).
@@
  • Extensive experiments demonstrate that our framework outperforms existing SVC methods on the benchmarks covering both autonomous driving and general scenes.
2 Related Work
2.1 Single-view Image and Video Compression
Single-view image and video compression eliminates spatial and temporal redundancies to achieve efficient encoding with minimal information loss. Traditional image codecs [7, 9, 51] rely on hand-crafted transforms (e.g., DCT) to remove spatial redundancy, while video codecs [59, 52, 9] use motion compensation to leverage content from previous frames. Despite their effectiveness, these manually designed approaches have fundamental limitations in optimizing compression efficiency.
Learning-based methods overcome these constraints by optimizing compression end-to-end. Ballé et al. [5] introduced a neural transform coding framework for image compression, significantly improving coding efficiency. Building on this, subsequent works [6, 47, 14, 23, 67, 62, 39, 32, 21, 22, 40, 41, 10] enhanced entropy models and network architectures, surpassing traditional codecs. In video compression, Lu et al. [43] pioneered the first neural video codec (DVC) by integrating learnable motion compensation into a neural image compression framework. Following this, research efforts [18, 4, 33, 27, 45, 24, 37, 28, 50, 34, 35, 36, 13, 12, 26] have progressively improved motion estimation, compensation mechanisms, and entropy models, leading to steady advancements in compression performance.
2.2 Stereo Image and Video Compression
Stereo image and video compression extends single-view methods by addressing cross-view redundancy through disparity compensation, which aligns stereo views to enhance reconstruction and compression efficiency. Traditional methods such as MV-HEVC [53] augment HEVC [52] with handcrafted disparity compensation but struggle to fully exploit cross-view dependencies.
Figure 1: The overall architecture of the proposed framework. At time step $t$, the input frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ is compressed by a series of modules, conditioned on the reconstructed frame pair $\{\bm{\hat{x}}_{t-1}^{L},\bm{\hat{x}}_{t-1}^{R}\}$ and feature pair $\{\bm{F}_{t-1}^{L},\bm{F}_{t-1}^{R}\}$ from the previous time step. In particular, the proposed HDC-FER and HDC-EM modules, which extend our hybrid disparity compensation (HDC) strategy to feature extraction and reconstruction, as well as entropy modeling, are integrated into both the motion and context compression components to reduce cross-view redundancy during the encoder–decoder transform and entropy prediction processes.
Recent advances in deep learning for stereo vision [49, 20, 29, 57, 65, 56, 15] have led to more flexible disparity compensation techniques. In stereo image compression, explicit methods leverage predefined disparity shifts or transformations for feature alignment [38, 17, 60], while implicit methods employ cross-attention for feature fusion [66, 61, 42], both of which significantly improve compression performance.
For stereo video compression, neural stereo video codecs (NSVCs) are also designed to eliminate cross-view redundancy. LSVC [11] sequentially compresses each view using offset-based disparity compensation, whereas LLSS [25] builds a cost volume with a Bishift operation, enabling parallel compression while capturing more comprehensive disparity information.
Broadly, existing disparity compensation techniques fall into two categories. Explicit alignment methods [38, 60, 25, 64, 17, 11] use predefined disparity shifts for feature warping, providing strong geometric guidance. However, they primarily focus on local pixel correspondences, making them less effective in occluded regions or when disparity estimates deviate from assumed shifts. In contrast, implicit alignment methods [31, 61, 66] employ cross-attention to perform global feature matching, which enhances robustness against occlusions and noise. Nevertheless, this approach comes with increased computational complexity and training challenges, often leading to slower convergence. By integrating both alignment strategies, our proposed hybrid disparity compensation (HDC) strategy combines the geometric reliability of explicit shifts with the adaptability of cross-attention, enabling more efficient and robust stereo video compression.
3 Methodology
3.1 Overall Architecture
Figure 2: The proposed hybrid disparity compensation module for feature extraction and reconstruction (HDC-FER). Given input features $\bm{K}^{L}$ and $\bm{K}^{R}$, 4D volumes $\bm{V}^{L}$ and $\bm{V}^{R}$ are generated by downsampling and shifting $\bm{K}^{L}$ and $\bm{K}^{R}$ [49, 29, 25], along the width dimension ($W$) from 1 to D. D is the max disparity value. An attention score $\bm{F}^{*}$ is then generated to weight and fuse the 4D volumes, producing aligned reference features $\bm{K}_{Ref}^{R}$ and $\bm{K}_{Ref}^{L}$ to enhance $\bm{K}^{R}$ and $\bm{K}^{L}$.
Given a stereo frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ at time step $t$, the goal of NSVC is to represent this frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ with fewer bits through an encoding process, with the original format can be restored via a reverse decoding process, while maintaining a specified level of reconstruction quality. In this paper, we propose an enhanced stereo video codec that builds upon modified video compression framework DCVC-TCM [50] to compress stereo video frames, an overview of our framework can be found in the left part of Fig. 1. With the input stereo frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ at time step $t$, motion vector pair $\{\bm{mv}_{t}^{L},\bm{mv}_{t}^{R}\}$ is initially estimated by the motion estimation network, using the reconstructed frame pair $\{\bm{\hat{x}}_{t-1}^{L},\bm{\hat{x}}_{t-1}^{R}\}$ from time step $t-1$. The estimated motion vector is then compressed by the motion compression network and reconstructed as $\{\bm{\hat{mv}}_{t}^{L},\bm{\hat{mv}}_{t}^{R}\}$. Subsequently, the motion compensation network generates multi-scale context features for each view using the feature pair $\{\bm{F}_{t-1}^{L},\bm{F}_{t-1}^{R}\}$ from the previous time step, and utilizes $\{\bm{\hat{mv}}_{t}^{L},\bm{\hat{mv}}_{t}^{R}\}$ to align the multi-scale context features to the current time step, producing an aligned multi-scale context feature pair $\{\bm{Ctx}^{L}_{1\sim 3},\bm{Ctx}^{R}_{1\sim 3}\}$. Following this, $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ is compressed by the context compression network, and reconstructed as $\{\bm{F}_{t}^{L},\bm{F}_{t}^{R}\}$, with $\{\bm{Ctx}^{L}_{1\sim 3},\bm{Ctx}^{R}_{1\sim 3}\}$ serving as conditional input. Finally, the reconstructed feature pair $\{\bm{F}_{t}^{L},\bm{F}_{t}^{R}\}$ is fed into the frame reconstruction network to obtain the reconstructed frame pair $\{\bm{\hat{x}}_{t}^{L},\bm{\hat{x}}_{t}^{R}\}$. Among the above five components, the motion estimation and motion compensation components mainly exploit temporal information to reduce transmission bit-stream, while the frame reconstruction component is a general module of reconstructing high quality image from processed feature. In this paper, we reduce cross-view redundancy in both the frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ and motion vector pair $\{\bm{mv}_{t}^{L},\bm{mv}_{t}^{R}\}$ by introducing the Hybrid Disparity Compensation (HDC) strategy. HDC is incorporated into both the Motion Compression and Context Compression modules to improve compression efficiency and reduce stereo video transmission overhead.
@@ 3.1 Overall Architecture
To achieve this, HDC integrates explicit pixel displacement, which provides a robust structural prior, with implicit cross-attention mechanisms for effective disparity compensation. As illustrated in the right part of Fig. 1, HDC first horizontally shifts features from both views to construct feature volumes at different disparity levels. A similarity map is then computed via an element-wise dot product between these volumes, followed by normalization to generate an attention score. This score is subsequently used in a cross-attention mechanism to adaptively align features across views, enabling efficient cross-view feature fusion.
Building on the HDC strategy, we introduce two HDC-based modules: Hybrid Disparity Compensation for Feature Extraction and Reconstruction (HDC-FER) and Hybrid Disparity Compensation for Entropy Modeling (HDC-EM). HDC-FER leverages the HDC mechanism to align and fuse cross-view features during feature extraction and reconstruction, thereby enhancing the quality of the recovered features. In contrast, HDC-EM incorporates aligned cross-view features as prior information during autoregressive entropy modeling to improve entropy prediction. Detailed descriptions of HDC-FER and HDC-EM are provided in Sec. 3.2.1 and Sec. 3.2.2.
3.2 Integration of HDC in our NSVC
3.2.1 HDC for Feature Extraction and Reconstruction
@@
In NSVCs, the encoder and decoder serve crucial roles in transforming the raw input and its latent representation into each other, effectively handling feature extraction and reconstruction. A sophisticated latent representation not only reconstructs the input with high fidelity but also remains amenable to compression, underscoring the importance of advanced encoder–decoder architectures in neural-based compression models. In this subsection, we explain how our Hybrid Disparity Compensation module for Feature Extraction and Reconstruction (HDC-FER) leverages cross-view information in features of both views to produce more informative features, thereby enhancing stereo video compression efficiency.
Figure 3: (a) The proposed Hybrid Disparity Compensation for Entropy Modeling (HDC-EM). The latent features $\bm{y}^{L}$ and $\bm{y}^{R}$, extracted by the encoder (see Fig. 1), are divided into $N$ equal slices along the channel dimension. These slices are processed autoregressively by the entropy estimation network $Est_{n}^{M}$ ($n\in\{1,\dots,N\}$, where $M$ denotes the left or right view) to generate coding symbols $\{\bm{\hat{y}}_{1}^{L},\dots,\bm{\hat{y}}_{N}^{L}\}$ and $\{\bm{\hat{y}}_{1}^{R},\dots,\bm{\hat{y}}_{N}^{R}\}$. Here, $\bm{Ctx}^{M}_{1\sim 3}$ represents the multi-scale context features described in Sec. 3.1; (b) The entropy estimation network $Est_{n}^{M}$ predicts the probability distribution for each latent slice $\bm{\hat{y}}_{n}^{M}$ using a context that comprises intra-view slices $\bm{\hat{y}}_{1\sim n-1}^{M}$, cross-view slices $\bm{\hat{y}}_{1\sim q}^{\overline{M}}$ (aligned to the current view via the HDC mechanism, with $q=n-1$ for the left view and $q=n$ for the right view), and spatial–temporal slices $\bm{\Phi}_{1\sim n}^{M}$. The coding symbol $\bm{\hat{y}}_{n}^{M}$ is then generated and encoded using the predicted mean $\bm{\mu}_{n}^{M}$ and scale $\bm{\sigma}_{n}^{M}$; (c) The spatial–temporal prior fusion network generates the context $\bm{\Phi}^{M}$ by fusing the hyperprior—extracted from $\bm{y}^{M}$ via the hyper encoder-decoder ($h_{a}$ and $h_{s}$)—with the multi-scale context features $\bm{Ctx}^{M}_{1\sim 3}$. The fused context is then evenly divided into $N$ slices to support entropy prediction. (Note: In the motion compression module, the context fusion network is omitted.)
@@
Our HDC-FER module is designed to enhance feature representation during feature extraction and reconstruction by leveraging HDC-based cross-view alignment. Specifically, HDC-FER performs bidirectional alignment, mapping features from each view to its counterpart, enabling mutual enhancement and improving stereo video compression efficiency. For simplicity, we denote the intermediate features of the left view and right view in the encoder-decoder as $\bm{K}^{L}$ and $\bm{K}^{R}$, respectively (see Fig. 2). Since there exists a significant disparity offset along the width dimension between $\bm{K}^{L}$ and $\bm{K}^{R}$, it is hard for the convolutional networks to aggregate cross-view information for the two views due to a limited receptive field. To address this, we follow prior stereo-processing works [29, 25, 49] and shift feature maps horizontally from 1 to a maximum disparity value, constructing 4D feature volumes $\bm{V}^{L}$ and $\bm{V}^{R}$ with the shape of $D\times C\times H\times W$:
@@
$\begin{split}\bm{K}^{L}_{\rm{Ref}}&=\operatorname{Conv3D}\left(\bm{F}^{*}\odot\bm{V}^{R}\right),\\ \bm{K}^{R}_{\rm{Ref}}&=\operatorname{Conv3D}\left(\bm{F}^{*}\odot\bm{V}^{L}\right),\\ \end{split}$ (3)
These reference features undergo upsampling and are added to the original inputs, enabling each view to enhance its feature representation using complementary information from the other view. Thus, HDC-FER integrates the HDC-based cross-attention mechanism into feature extraction and reconstruction, ensuring that each view benefits from the most relevant cross-view features. This results in a more informative and compact latent representation, improving stereo video compression efficiency.
3.2.2 HDC for Cross-View Entropy Modeling

### v2

@@ Abstract
I Introduction
As stereoscopic cameras become increasingly common in newly launched smart products, such as virtual reality headsets and autonomous vehicles, the demand for efficient storage and transmission of large volumes of stereo video is growing. Stereo video compression (SVC) [1, 2, 3, 4] has recently gained significant attention for its ability to encode stereo videos into compact bitstreams, reducing both storage and transmission costs. However, developing an effective SVC algorithm remains a non-trivial research problem. A key challenge is reducing cross-view redundancy, where the disparity compensation is central to this effort. By aligning stereo views, disparity compensation enables the extraction of missing details and the reuse of shared information, which enhances reconstruction quality and improves compression efficiency. Traditional SVC [2, 1] rely on handcrafted disparity compensation methods, but these approaches are limited by the constraints of manually designed structures. The recent success of deep neural networks (DNNs) in stereo visual data processing [5, 6, 7, 8, 9, 10, 11] has paved the way for neural-based disparity compensation mechanisms. These modern techniques offer a more flexible and accurate solution for reducing cross-view redundancy in SVC.
@@ I Introduction
We conduct extensive experiments to evaluate our framework on widely used automotive stereo video benchmarks, including KITTI 2012 [19] and KITTI 2015 [20]. In contrast to previous NSVC methods [4, 3]—which have been evaluated solely on automotive scenes—we also test our approach on general multi-view scenes using the Nagoya dataset [21], a standard stereo video benchmark in MV-HEVC [1, 2]. Experimental results highlight the effectiveness of our framework: it surpasses both neural and traditional SVC methods in automotive scenarios and achieves performance comparable to MV-HEVC in general multi-view scenes. Our contributions can be summarized as follows:
@@
  • We propose a hybrid disparity compensation, i.e., HDC, strategy that leverages explicit pixel displacement as a robust disparity prior and incorporates implicit cross-attention for improved compensation performance.
@@
  • We develop an end-to-end optimized NSVC framework by integrating our HDC into key components, including cross-view feature extraction and reconstruction and cross-view entropy modeling.
@@
  • Extensive experiments demonstrate that our framework outperforms existing SVC methods on autonomous driving benchmarks and achieves performance comparable to the standard MV-HEVC in general multi-view scenes.
II Related Work
II-A Video compression
Traditional video codecs [22, 23, 24] employ hand-crafted motion compensation modules to exploit temporal dependencies across frames. Despite their proven effectiveness, these hand-crafted modules face inherent limitations in optimizing the rate-distortion trade-off.
Recently, NVC has seen rapid advancements. Mainstream frameworks can be broadly categorized into residual coding [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36] and conditional coding paradigms [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. Specifically, residual coding methods explicitly subtract predicted frames from the current frame. For instance, Lu et al. [51] introduced the first end-to-end NVC framework by incorporating learnable motion estimation and compensation. These schemes depend strongly on precise spatial alignment, as it is critical for effective explicit motion compensation. On the other hand, another category conditional coding, captures temporal dependencies directly within the latent feature space, enabling more efficient and adaptable compression. In particular, the DCVC family [37, 39, 41, 43, 45, 50] adopts advanced conditional coding architectures and has demonstrated significant improvements in video compression performance across a range of benchmarks.
While these methods achieve impressive results in single-view video compression, they are not readily applicable to MVC task due to their lack of consideration for cross-view redundancy inherent in multi-view settings, thereby limiting potential performance gains. To address this limitation, we propose a novel disparity compensation module and build upon DCVC-TCM [50], a state-of-the-art yet relatively lightweight member of the DCVC family, to strike a balance between compression performance and efficiency.
II-B Stereo Image Compression
Stereo image compression (SIC) techniques typically build upon conventional image compression methods by additionally leveraging cross-view redundancy through disparity compensation. This compensation process aligns stereo image pairs to improve both reconstruction quality and coding efficiency. Traditional SIC algorithms implement the disparity mechanism using hand-crafted modules such as non-separable vector lifting [52], wavelet [53], and variable-block [54]. However, similar to traditional video compression methods, such hand-crafted modules are inherently limited in their ability to support end-to-end optimization.
Recently, neural SIC approaches have achieved tremendous success by employing deep learning-based disparity compensation strategies, which can be broadly categorized into two types: explicit and implicit alignment. Specifically, explicit alignment methods [12, 14, 15, 55] utilize predefined disparity shifts to warp features across views, leveraging strong geometric priors such as DSIC [12], HESIC [15], DispSIC [14], and MASIC [55]. For example, Deng et al. [15] applies homography-based warping to align stereo features based on estimated planar transformations. Nevertheless, these methods rely heavily on accurate local correspondences, making them vulnerable to occlusions and disparity estimation errors.
On the other hand, implicit alignment methods [16, 17, 18, 13, 16] perform global feature interaction using cross-attention mechanisms to generate implicit similarity maps. For example, SASIC [13] and ECSIC [17] adopt attention-based matching strategy that enhances robustness and effectively handles occlusions and noise. However, generating such implicit similarity maps is computationally intensive during both training and inference, leading to slower convergence and increased training costs.
Although both categories have achieved notable success in the SIC domain, extending those SIC strategies to SVC presents a non-trivial challenge. In SVC, both performance and efficiency are critical, as disparity compensation must be applied to every stereo image pair at each time step. Poor performance or high computational cost can lead to significant error accumulation and increased latency. To address this, we propose a hybrid compensation strategy that leverages the strengths of both paradigms by combining the simplicity and robustness of explicit alignment with the high representational capacity of implicit cross-attention mechanisms. This hybrid design enables efficient optimization while maintaining high disparity compensation accuracy across time steps.
Figure 1: Left: Overview of our Neural Stereo Video Compression (NSVC) framework. At time step $t$, the input stereo frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ is compressed and reconstructed into $\{\bm{\hat{x}}_{t}^{L},\bm{\hat{x}}_{t}^{R}\}$ using five key coding components. This process is conditioned on the previously reconstructed frame pair $\{\bm{\hat{x}}_{t-1}^{L},\bm{\hat{x}}_{t-1}^{R}\}$ and the corresponding feature pair $\{\bm{F}_{t-1}^{L},\bm{F}_{t-1}^{R}\}$ from time step t1t-1 as in established neural video compression (NVC) methods [50]. Middle: In the two key components, Motion Compression and Context Compression, we apply and extend our Hybrid Disparity Compensation (HDC) strategy. Specifically, HDC-FER mechanisms are used within their encoders and decoders to enhance the intermediate feature pair $\{\bm{K}_{t}^{L},\bm{K}_{t}^{R}\}$, while HDC-EM mechanisms are used within their entropy model to improve the quantized latent feature pair $\{\bm{\hat{y}}_{t}^{L},\bm{\hat{y}}_{t}^{R}\}$. Right: High-level syntax of the proposed HDC strategy. We first apply explicit pixel displacement to construct a similarity map, which is then used in a cross-attention mechanism to implicitly align features across views.
II-C Multi-view Video Compression
MVC techniques aim to compress video sequences captured from multiple synchronized viewpoints. Unlike single-view video coding, MVC simultaneously leverages temporal redundancy within each view and inter-view redundancy across different views through joint motion and disparity compensation. For instance, established standards such as MVC [1] and MV-HEVC [2] extend the conventional H.264 [22] and H.265 [23] architectures by incorporating inter-view reference frames and hierarchical prediction structures.
SVC serves as a foundational technique for those MVC frameworks. To enhance SVC performance, two NSVC approaches have emerged, leveraging learning-based disparity compensation techniques. Specifically, LSVC [4] employs a sequential compression scheme with deformable-based disparity warping, while LLSS [3] introduces a novel Bishift operation to construct cost volumes for parallel disparity modeling. To achieve high efficiency, both methods adopt explicit disparity shift-based mechanisms. However, as discussed in Sec. II-B, although explicit disparity compensation is computationally efficient, its performance is often constrained by limited robustness to errors and occlusions. In this work, we propose a hybrid compensation strategy that balances efficiency and accuracy, enabling high-performance SVC.
III Methodology
III-A Overall Architecture
Given a stereo frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ at time step $t$, the objective of NSVC is to represent this pair using fewer bits via an encoding process, such that the original data can be accurately reconstructed through a corresponding decoding process, while preserving a specified level of visual quality. In this work, we present an enhanced stereo video codec built upon a modified version of the DCVC-TCM video compression framework [50], which is adapted to efficiently compress stereo video frames. An overview of the proposed framework is illustrated in the left part of Fig. 1.
Motion Estimation. Given the input stereo frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ at time step $t$ and the reconstructed stereo frame pair $\{\bm{\hat{x}}_{t-1}^{L},\bm{\hat{x}}_{t-1}^{R}\}$ from the previous time step (retrieved from the decoded buffer), the motion estimation network initially estimates the motion vector pair $\{\bm{mv}_{t}^{L},\bm{mv}_{t}^{R}\}$.
Motion Compression. The estimated motion vector pair $\{\bm{mv}_{t}^{L},\bm{mv}_{t}^{R}\}$ is subsequently compressed by the motion compression network and reconstructed as $\{\bm{\hat{mv}}_{t}^{L},\bm{\hat{mv}}_{t}^{R}\}$. The motion compression network consists of an encoder, an entropy model, and a decoder, which are responsible for feature extraction, entropy modeling, and feature reconstruction, respectively. The encoder transforms the input motion vectors into a latent feature pair, which is then quantized and losslessly compressed using arithmetic coding. The entropy model estimates the probability distribution of the quantized latent feature pair to facilitate cross-entropy coding. The decoder reconstructs the motion vectors from the quantized latent features. During the feature extraction (resp., reconstruction) process, an intermediate feature pair is produced after each downsampling (resp., upsampling) operation.
Motion Compensation. The motion compensation network generates multi-scale context features for each view using the feature pair $\{\bm{F}_{t-1}^{L},\bm{F}_{t-1}^{R}\}$ from the previous time step, retrieved from the decoded buffer. It then utilizes the motion vector pair $\{\bm{\hat{mv}}_{t}^{L},\bm{\hat{mv}}_{t}^{R}\}$ to align the multi-scale context features to the current time step, producing an aligned multi-scale context feature pair $\{\bm{Ctx}^{L}_{1\sim 3,t},\bm{Ctx}^{R}_{1\sim 3,t}\}$.
Context Compression. The stereo frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ is compressed by the context compression network and reconstructed as $\{\bm{F}_{t}^{L},\bm{F}_{t}^{R}\}$, with the multi-scale context feature pair $\{\bm{Ctx}^{L}_{1\sim 3,t},\bm{Ctx}^{R}_{1\sim 3,t}\}$ serving as conditional input. This context feature pair is generated by a multi-scale temporal context mining network and incorporated into the feature extraction (resp., reconstruction) process to effectively exploit temporal information, following the design in [50]. Similar to the motion compression network, the context compression network also consists of an encoder, an entropy model, and a decoder, which together generate the latent feature pair for compression and the intermediate feature pair at different scales during feature extraction and reconstruction process.
Frame Reconstruction. The reconstructed feature pair $\{\bm{F}_{t}^{L},\bm{F}_{t}^{R}\}$ is fed into the frame reconstruction network to generate the reconstructed frame pair $\{\bm{\hat{x}}_{t}^{L},\bm{\hat{x}}_{t}^{R}\}$ at the current time step.
Among the five components described above, the motion estimation and motion compensation modules primarily leverage temporal information to reduce the transmission bitstream, whereas the frame reconstruction module serves as a general component for recovering high-quality images from the processed features. To further address the cross-view redundancy inherent in stereo videos, this work aims to enhance the quality of the quantized latent feature pair and intermediate feature pair through a carefully designed disparity compensation mechanism.
III-B Hybrid Disparity Compensation
As mentioned earlier, our goal is to improve the quality of intermediate and quantized latent features for both the left and right views during the motion and context compression processes. To this end, we exploit cross-view correlations to enhance feature representations. Specifically, in both the motion and context compression modules, we align the intermediate or latent features of the left (resp., right) view with those from the right (resp., left) view.
To enable effective cross-view feature alignment, we employ our Hybrid Disparity Compensation (i.e., HDC) strategy. HDC is integrated into three key components of each module: the multi-scale encoder, the entropy model, and the multi-scale decoder, as illustrated in the central part of Fig. 1111For simplicity, we denote the multi-scale intermediate feature pairs produced by the encoder and decoder as $\{K_{t}^{L},K_{t}^{R}\}$, and the latent feature pairs as $\{y_{t}^{L},y_{t}^{R}\}$, for both the motion and context compression modules.. First, to produce higher-quality intermediate feature pairs $\{K_{t}^{L},K_{t}^{R}\}$, we propose Hybrid Disparity Compensation for Feature Extraction and Reconstruction (i.e., HDC-FER). This module aligns and fuses cross-view features during both the feature extraction and reconstruction stages, thereby enhancing the quality of the resulting features. Second, to improve the latent feature pairs $\{y_{t}^{L},y_{t}^{R}\}$ for more effective entropy modeling in both views, we introduce Hybrid Disparity Compensation for Entropy Modeling (i.e., HDC-EM). This technique aligns cross-view features and leverages them as prior information to enhance the accuracy of entropy estimation.
@@ III-B Hybrid Disparity Compensation
In general, both HDC-FER and HDC-EM are specialized instantiations of our high-level HDC method, as illustrated in the right part of Fig. 1 Specifically, HDC combines explicit pixel displacement, providing a robust structural prior with implicit cross-attention mechanisms for effective disparity compensation. To capture explicit pixel displacement, we horizontally shift features from both views to construct feature volumes across multiple disparity levels. A similarity map is then computed using an element-wise dot product between these volumes, followed by normalization to produce an attention score. This score is subsequently utilized in a cross-attention mechanism to implicitly align features across views. Detailed descriptions of HDC-FER and HDC-EM are provided in Sec.III-C1 and Sec. III-C2, respectively.
III-C Hybrid Disparity Compensation for NSVC
III-C1 HDC for Feature Extraction and Reconstruction (HDC-FER)
@@
In NSVCs, the encoder and decoder serve crucial roles in transforming the raw input and its latent representation into each other, effectively handling feature extraction and reconstruction. A sophisticated latent representation not only reconstructs the input with high fidelity but also remains amenable to compression, underscoring the importance of advanced encoder–decoder architectures in neural-based compression models. In this subsection, we explain how our Hybrid Disparity Compensation module for Feature Extraction and Reconstruction (HDC-FER) leverages cross-view information in features of both views to reduce cross-view redundancy during encoding and produce more informative features during decoding, thereby enhancing stereo video compression efficiency.
@@
Our HDC-FER module is designed to enhance feature representation during feature extraction and reconstruction by leveraging HDC-based cross-view alignment. Specifically, HDC-FER performs bidirectional alignment, mapping features from each view to its counterpart, enabling mutual enhancement and improving stereo video compression efficiency. For simplicity, we denote the intermediate features of the left view and right view in the encoder-decoder as $\bm{K}^{L}$ [^2] For brevity, the time index $t$ is omitted from all variables in Sec. III-C, including variables shown in figures.. and $\bm{K}^{R}$, respectively (see Fig. 2). Since there exists a significant disparity offset along the width dimension between $\bm{K}^{L}$ and $\bm{K}^{R}$, it is hard for the convolutional networks to aggregate cross-view information for the two views due to a limited receptive field. To address this, we follow prior stereo-processing works [7, 3, 5] and shift feature maps horizontally from 1 to a maximum disparity value, constructing 4D feature volumes $\bm{V}^{L}$ and $\bm{V}^{R}$ with the shape of $D\times C\times H\times W$:
@@
Figure 2: The proposed Hybrid Disparity Compensation Module for Feature Extraction and Reconstruction (HDC-FER). We input the intermediate feature pair $\{\bm{K}^{L},\bm{K}^{R}\}$ produced by the encoder and decoder modules from Motion and Context Compression (See Fig. 1). We first perform downsampling and horizontally shifting operation for the intermediate feature $\bm{K}^{L}$ (resp., $\bm{K}^{R}$) to obtain the 4D disparity volume $\bm{V}^{L}$ (resp., $\bm{V}^{R}$), where we shift the $\bm{K}^{L}$ (resp., $\bm{K}^{R}$) across a disparity range from 11 to $D$ and $D$ denotes the maximum disparity value. Subsequently, an attention score $\bm{F}^{*}$ (i.e., similarity map) is generated by applying standard operations, element-wise dot product, $\operatorname{Softplus}$ function and $\operatorname{Tanh}$ function. Then, we apply this score into the previously produced cost-volume $\bm{V}^{L}$ (resp., $\bm{V}^{R}$). After a refinement, we can obtain final aligned reference feature $\bm{K}_{Ref}^{R}$ (resp., $\bm{K}_{Ref}^{L}$), which will be added back to the intermediate feature $\bm{K}^{L}$ (resp., $\bm{K}^{R}$) to produce new intermediate feature $\bm{K}_{new}^{L}$ (resp., $\bm{K}_{new}^{L}$).
@@
[…]
These reference features undergo upsampling and are added to the original inputs, enabling each view to enhance its feature representation using complementary information from the other view. Thus, HDC-FER integrates the HDC-based cross-attention mechanism into feature extraction and reconstruction, ensuring that each view benefits from the most relevant cross-view features. During encoding, it helps both views reduce cross-view redundancy, thereby improving compression efficiency. During decoding, it yields a more informative and compact latent representation that further improves the reconstruction quality of stereo video frames.
III-C2 HDC for Cross-View Entropy Modeling (HDC-EM)